# Hiba Arnaout1, Noy Sternlicht2, Tom Hope2,3, Iryna Gurevych1

### v1

@@ Abstract
Understanding the impact of scientific publications is crucial for identifying breakthroughs and guiding future research. Traditional metrics based on citation counts often miss the nuanced ways a paper contributes to its field. In this work, we propose a new task: generating nuanced, expressive, and time-aware impact summaries that capture both praise (confirmation citations) and critique (correction citations) through the evolution of fine-grained citation intents. We introduce an evaluation framework tailored to this task, showing moderate to strong human correlation on subjective metrics such as insightfulness. Expert feedback from professors reveals a strong interest in these summaries and suggests future improvements. Data and code are made available.  [^1] https://ukplab.github.io/arxiv2025-generating-impact-summaries
@@ 1 Introduction
We break the task of generating impact summaries for research papers (Section 2.1) into two subtasks. First, given all the citation contexts of a paper, we identify the “impact revealing” citations that directly interact with the paper, and their specific intents, whether in confirmation of its ideas (e.g., “method use”) or in correction (e.g., “identifying limitations and proposing refinements”). Second, using only the impact-revealing citation contexts, their dates and intents, we generate an impact summary that captures the paper’s influence over time. To the best of our knowledge, this is the first attempt to express scientific impact through time-aware textual summaries derived from citation context analysis. An overview is shown in Figure 1. To implement these subtasks, our method, in Section 2, uses in-context learning (ICL) with Large Language Models (LLMs) to generate fine-grained citation intents in the first stage. In the second stage, it filters for impact-revealing citation contexts based on the generated intents and feeds them to an LLM to generate semi-structured impact summaries under a pre-defined schema.
Existing work on scientific impact analysis predominantly focuses on citation and similar numerical counts as the main indicator of impact Gu and Krenn (2024), ignoring the exact reasons a paper has been cited, which can often be “unimpactful”, e.g., citing a paper to acknowledge background work on a topic. Work on citation context analysis Lauscher et al. (2022) studies the role citations play, but is based on coarse categories and is not directly intended for impact analysis. The goal of our new task is to create a narrative about a paper’s impact using fine-grained citation intent analysis. Moreover, existing work typically emphasizes praise (confirmatory citations)  Zhang et al. (2024); Valenzuela et al. (2015), but scientific progress relies on both confirmation and correction Catalini et al. (2015), as ideas advance through adoption as well as critique and refinement. In this work, we ensure that our generated impact summaries capture both aspects of impact (when applicable), as shown in Figure 1. To support this broader view of impact, we extend the existing PST-Bench data set Zhang et al. (2024), which only focuses on “motivation” and “inspiration” intentions to consider an impact revealing citation, by introducing, in Section 2.5, a new ground-truth data set of 4000 citation contexts labeled for whether they reveal impact. Our extended dataset also covers not only additional intents of confirmatory citations (e.g., method use), but also correction citations (e.g., method refinement). For detailed comparisons with related work, see Appendix A and Section 5.
@@ 2.1 Task definition
Citation intent refers to the purpose (the “why”) behind citing paper $p$ in paper $p^{\prime}$, normally inferred from the citation context.
@@
Figure 1 shows the intent “acknowledge background work”, for the context that simply refers to prior work in a general manner, without actively building upon or challenging it. Hence, this is not considered impact-revealing. In contrast, the intent “method used”, for the context describing the implementation of a popular approach is indeed impact-revealing.
@@
2.2 Identifying impact-revealing citation intents
We use ICL with LLMs to generate fine-grained intents and classify them as either “impact-revealing” or “other”. To facilitate this, we manually craft a prompt that incorporates a set of examples covering various citation intents, and ensuring the model can learn from them to generate fine-grained intents and classify them correctly (prompt and examples are in Appendix B and  E.1, respectively). This prompt is designed to capture both the confirmatory and correction citations, allowing for a nuanced understanding of how the cited paper influences the citing paper. It aims to generate concise but expressive labels that clearly explain the reason behind the citation without being overly general or overly specific, e.g., “identifying knowledge gap in literature” (more examples in Appendix E.1). To examine the quality of our method, we conduct experiments in Sections 3.1 and  3.2.
@@ 2.4 Evaluation metrics
A key challenge in evaluating our summaries is the lack of gold-standard datasets, due to the task’s novelty. We attempted to construct gold summaries using sources like Wikipedia articles on influential papers and Test of Time Award webpages, but these proved inadequate. Such sources cover only a few papers and offer details about the content of the paper itself [^2] E.g., Wikipedia article for Vaswani et al. (2017): https://en.wikipedia.org/wiki/Attention_Is_All_You_Need. Similarly, Test of Time award announcements [^3] E.g., WSDM 2022 Test of Time Award Winner:https://www.wsdm-conference.org/2022/timetable/event/wsdm-awards-program-test-of-time-presentation/ focus on broad praise, citation counts, and author accomplishments, lacking usage-based insights. Web searches using queries like “what is the scientific impact of paper X” mostly returned different versions of the paper itself, or occasionally, blog posts highlighting novelty, or explaining the paper in simpler terms. Given these limitations, we introduce a new, automated, reference-free evaluation framework with metrics capturing both the trustworthiness and informativeness of the summaries, and later demonstrate its moderate to strong correlation with human judgment (Appendix G.3).
@@
Trustworthiness. Measuring summary correctness is challenging due to the large gap between input and output lengths, compressing thousands of citation contexts into a few sentences. Without a gold reference, determining what content should be included or excluded becomes difficult. Additionally, time-augmented impact descriptions must be accurate both in content and assigned time period. Inspired by evaluation frameworks in the RAG setting Ru et al. (2024); Shahul et al. (2023); Asai et al. (2024), which aligns with our query-based summarization approach, we define:
(1) Faithfulness: Our time-Aware faithfulness metric examines whether details in the impact summary, i.e., the impact description elaborating on the dominant intent of an impact period, is grounded in the paper’s citation contexts (i.e., the LLM’s input context). We first split the impact summaries into impact descriptions representing different time periods (as defined in the output schema in Appendix C). Next, we instruct an LLM to verify each against citations from the corresponding period. Note that a summary impact description is verified against many citations at once and not in a simple pairwise entailment, since a single impact description can encompass information from multiple citations or discuss trends, which are only possible to verify by inspecting many sources (e.g., “the paper is frequently used for…”). The evaluator LLM is instructed to assess the impact description against the provided citations, determine its faithfulness, and justify the decision with a list of supporting citations. This procedure resembles the task of a human annotator evaluating the faithfulness of a machine-generated text Kim et al. (2024).
(2) Coverage: Coverage is defined as the ratio of impact-revealing intents mentioned in the summary. For example, if a paper is cited for its method use, for inspiring new research, and for exposing limitations and refinement, the summary should reflect all these intents, including relevant details. Inspired by the evaluation rubrics for coverage proposed in Asai et al. (2024), which defines it as the topics mentioned (i.e., the diversity of impact-revealing intents in our case), we develop a two-step evaluation. In the first step, we cluster the list of fine-grained intents under similar topics, e.g., “method used in the legal domain”. These labels might have slightly different granularity, which depends on how much details the members (i.e., intents) of that cluster offer. Next, the evaluator LLM takes both the impact summary and the list of cluster titles that this summary should ideally cover, and returns a list of topics that were actually covered. Finally, we divide the number of intents mentioned in the summary by the total number of impact-revealing intent clusters. Prompts for faithfulness and coverage are in Appendix D.
(3) Citation Year Compliance: We implement this metric as a script that flags citations falling outside the target impact period.
Informativeness. We also assess the informativeness of impact summaries, as trustworthy ones may still be unhelpful if they lack meaningful insights. Grounded in our definitions (Section 2.1), the informativeness of an impact summary depends on its ability to clearly describe the paper’s direct influence on other papers, track the types of influence (intents) over time, and provide details inferred from the citation contexts. We define the following metrics:
(1) Insightfulness measures how well the impact summary articulates the paper’s direct influence on other works.
(2) Trend Awareness evaluates the extent to which the impact summary identifies and distinguishes between different periods of the paper’s influence over time.
(3) Specificity assesses whether the impact summary includes concrete examples, such as techniques, influenced by the paper. To systematically evaluate these criteria, we leverage G-Eval Liu et al. (2023c), a framework that uses LLMs with chain-of-thought (CoT) reasoning to assess the impact summaries. By integrating structured evaluation steps, G-Eval enables LLMs to reason step-by-step, improving the reliability and depth of assessments. We transform our criteria into prompts in Appendix D.
To construct a new dataset for classifying citation contexts as “impact-revealing” or not, we start by building on the PST-Bench dataset Zhang et al. (2024), which focuses on positive influence through “inspiration” and “motivation”. We augment this by first manually crafting textual patterns for both confirmation and correction intents, and use GPT-4 to generate variations. After crawling 200k random citation contexts (from the Semantic Scholar Academic Graph (S2AG) [^4] https://www.semanticscholar.org/product/api) and checking if they match any of the textual patterns, we randomly sample 2k instances of impact-revealing contexts (50% for confirmation and correction citations each). Additionally, we sample 2k non-impact-revealing citations (“other”) directly from the PST-Bench dataset (references that were not annotated as influential). We also ensure none of these match any of our impact-revealing citation patterns. This results in a balanced 4k citation context dataset, which we release alongside this work. Details on the construction and textual patterns are in Appendix F.1.
3 Experiments
3.1 In-context-learning for fine-grained intent generation
Zero-shot vs. ICL. To evaluate LLMs’ ability to generate fine-grained intents and classify them as “impact-revealing” or “non-impact-revealing”, we manually select 10 citation contexts (Appendix E.1). Our selection criteria was based on diversity, a set that contains impact-revealing (praise, critique) and non-impact-revealing (“other”) intents. The “other” class includes incidental mentions or contexts lacking sufficient information to understand the reason behind it. We use our prompt (Appendix B) with and without these examples to generate and classify intents, using GPT4o-mini, of 200 randomly sampled citation contexts. Two annotators (a postdoc and a PhD student) assess: (i) whether the classification as impact-revealing is correct, (ii) whether the generated intent accurately reflects the citation reason, and (iii) whether the intent is concise, as LLM results can be highly verbose.
Results, including per-field metrics and qualitative examples, are in Appendix E.2. We find that including examples significantly improves performance across all metrics, with gains of 29%, 55%, and 39% in precision, recall, and F1, respectively. Inter-annotator agreement (F1 agreement) is 0.86 for the impact-revealing classifier, 0.77 and 0.85 for the correctness and conciseness of the generated fine-grained intents, respectively.
Effect of different number of shots (K). We conduct an experiment to test our ICL prompt with different Ks. We manually polished the 200 instances from the previous experiment and randomly split the set into train 40%, dev 30%, test 30%. For the number of shots, we test from K=1 to 80. For every dev and test instance, we run the prompt 5 times and report the average, using GPT4o-mini, shuffling the order and selection of the shots. Results are plotted in Figure 2. Increasing the number of shots has an effect on all metrics, outperforming both the ZS (horizontal dashed lines), and the always predict impact-revealing baselines (dotted lines). Performance improves sharply up to K = 50, after which gains are minimal. At K = 50, metrics plateau with precision at 0.90, recall at 0.94, F1 at 0.92, and accuracy at 0.92.
Figure 2: Effect of different number of shots (K) on detecting impact-revealing citations. Zero-shot: horizontal dashed lines; always predict impact-revealing baseline: dotted lines. This shows that adding a reasonable number of examples (around 50) is sufficient to yield a 60% improvement in recall.
Cited Papers Citation Intents (%)
PS MD CS
All 6535 5347 3466
Recent 6535 5743 5842
[…]
Highly cited 6634 3664 3862
[…]
How does citation intent vary across fields? Table 2 analyzes citation intents across 70k citation contexts from psychology, medicine, and computer science papers. It shows that, generally, psychology citations tend to lean towards impact-revealing, computer science towards the non-impact-revealing (“other”) side, with the exception of more recent papers, and medicine has a more balanced split. Looking at specific examples, we notice that citations in psychology papers often have a stronger subjective tone, particularly in correction citations. Moreover, in recent years, computer science shows a shift toward more impact-revealing citations, likely reflecting the novelty and immediate impact of AI research before entering the "legacy" phase with more background-oriented citations. Details and examples are in Appendix F.
@@ 3.2 Can existing intent classifiers detect impact-revealing citations?
Comparison with existing methods. We compare our impact-revealing classifier against the following popular intent classification methods: (i) Meaningful Citations Valenzuela et al. (2015): a supervised machine learning method to classify citation intents into meaningful or non-meaningful, by leveraging context and citation metadata; (ii) Structural Scaffolds Cohan et al. (2019): a multi-task model which enhances the main task by integrating auxiliary tasks; (iii) Multi-cite Lauscher et al. (2022): a multi-label classification method that predicts multiple intents using a fine-tuned SciBERT Beltagy et al. (2019). We map their intent classes to either “impact-revealing” or “other”. More details are provided in Appendix F.2. For our method, we use the ICL prompt, with K=50 and LLM=GPT4o-mini. For every test instance, we run our prompt 3 times (shuffling the order of the shots), and take a majority vote for the classification. We note that the results of the 3 runs have a full agreement of 72%.
Results. Numerical results are reported in Table 1. Our method demonstrates the best performance on this task, with the most significant improvement in recall, outperforming the second-best method by 19 percentage points. This is particularly crucial for our goal in generating impact summaries from impact-revealing citations because the highest recall in this context means that our method ensures that fewer meaningful citations are missed. The qualitative results are shown in the appendix F.2.

### v2

@@ Abstract
Understanding the impact of scientific publications is crucial for identifying breakthroughs and guiding future research. Traditional metrics based on citation counts often miss the nuanced ways a paper contributes to its field. In this work, we propose a new task: generating nuanced, expressive, and time-aware impact summaries that capture both praise (confirmation citations) and critique (correction citations) through the evolution of fine-grained citation intents. We introduce an evaluation framework tailored to this task, showing moderate to strong human correlation on subjective metrics such as insightfulness. Expert feedback from professors reveals a strong interest in these summaries and suggests future improvements. Data and code are made available.  [^1] https://ukplab.github.io/acl2026-generating-impact-summaries
@@ 1 Introduction
We break the task of generating impact summaries for research papers into two subtasks. First, given all the citation contexts of a paper, we identify the “impact revealing” citations that directly interact with the paper, and their specific intents, whether in confirmation of its ideas (e.g., “method use”) or in correction (e.g., “identifying limitations and proposing refinements”). Second, using only the impact-revealing citation contexts, their dates and intents, we generate an impact summary that captures the paper’s influence over time. To the best of our knowledge, this is the first attempt to express scientific impact through time-aware textual summaries derived from citation context analysis. An overview is shown in Figure 1. To implement these subtasks, our method, in Section 2, uses in-context learning (ICL) with Large Language Models (LLMs) to generate fine-grained citation intents in the first stage. In the second stage, it filters for impact-revealing citation contexts based on the generated intents and feeds them to an LLM to generate semi-structured impact summaries under a pre-defined schema.
Existing work on scientific impact analysis predominantly focuses on citation and similar numerical counts as the main indicator of impact Gu and Krenn (2024), ignoring the exact reasons a paper has been cited, which can often be “unimpactful”, e.g., citing a paper to acknowledge background work on a topic. Work on citation context analysis Lauscher et al. (2022) studies the role citations play, but is based on coarse categories and is not directly intended for impact analysis. The goal of our new task is to create a narrative about a paper’s impact using fine-grained citation intent analysis. Moreover, existing work typically emphasizes praise (confirmatory citations)  Zhang et al. (2024a); Valenzuela et al. (2015), but scientific progress relies on both confirmation and correction Catalini et al. (2015), as ideas advance through adoption as well as critique and refinement. In this work, we ensure that our generated impact summaries capture both aspects of impact (when applicable), as shown in Figure 1. To support this broader view of impact, we extend the existing PST-Bench data set Zhang et al. (2024a), which only focuses on “motivation” and “inspiration” intentions to consider an impact revealing citation, by introducing, in Section 2.5, a new ground-truth data set of 4000 citation contexts labeled for whether they reveal impact. Our extended dataset also covers not only additional intents of confirmatory citations (e.g., method use), but also correction citations (e.g., method refinement). For detailed comparisons with related work, see Appendix A and Section 5.
@@ 2.1 Definitions
Fine-grained citation intent refers to the specific, nuanced reason (the “why”) behind citing paper $p$ in paper $p^{\prime}$, as inferred from the citation context. Unlike predefined or categorical intent schemes, fine-grained citation intent is expressed in free-form text and does not rely on a fixed set of labels or a predefined taxonomy.
A few examples are in Table 1.
@@
Figure 1 shows the intent “acknowledge background work”, for the context that simply refers to prior work in a general manner, without actively building upon it. Hence, this is not considered impact-revealing. In contrast, the intent “method used”, describing the implementation of a popular approach is indeed impact-revealing.
@@
2.2 Identifying impact-revealing intents
We use ICL with LLMs to generate fine-grained intents and classify them as either “impact-revealing” or “other”. To facilitate this, we manually craft a prompt that incorporates a set of examples covering various citation intents, and ensuring the model can learn from them to generate fine-grained intents and classify them correctly (prompt and examples are in Appendix B and Table 1, respectively). This prompt is designed to capture both the confirmatory and correction citations, allowing for a nuanced understanding of how the cited paper influences the citing paper. It aims to generate concise but expressive labels that clearly explain the reason behind the citation without being overly general or overly specific, e.g., “identifying knowledge gap in literature” (more examples in Appendix E.1). To examine the quality of our method, we conduct experiments in Sections 3.1 and  3.2.
Citation context Intent Class
..we apply a minimization process [1]. use of minimization methodology impact-revealing
Quirk and Poon (2017) and Peng et al. (2017) build two distantly supervised datasets without human annotation, which may make the evaluation less reliable.. we present .. a large-scale human-annotated document-level RE dataset.. criticizing existing datasets and proposing a better one impact-revealing
Table 1: Training examples for fine-grained intent generation and classification. Full table in Appendix E.1.
@@ 2.4 Evaluation metrics
Evaluating our summaries is challenging due to the lack of gold-standard datasets for this novel task. We explored sources like Wikipedia articles on influential papers  [^2] E.g., Article for Vaswani et al. (2017): https://en.wikipedia.org/wiki/Attention_Is_All_You_Need and Test of Time Award pages [^3] E.g., WSDM 2022 Test of Time Award Winner: https://www.wsdm-conference.org/2022/timetable/event/wsdm-awards-program-test-of-time-presentation/, but these were insufficient, as they cover few papers and focus on content or general praise rather than usage-based insights. Web searches (e.g., “what is the impact of paper X”) mostly returned the original paper or explanatory blog posts. Given these limitations, we introduce a new automated, reference-free evaluation framework that measures trustworthiness and informativeness, and correlates moderately to strongly with human judgment (Appendix G.4).
Trustworthiness. Measuring summary correctness is challenging due to the large gap between input and output lengths, compressing thousands of citation contexts into a few sentences. Without a gold reference, determining what content should be included or excluded becomes difficult. Additionally, time-augmented impact descriptions must be accurate both in content and assigned time period. Inspired by evaluation frameworks in the RAG setting Ru et al. (2024); ES et al. (2024); Asai et al. (2024a), which aligns with our query-based summarization approach, we define: (1) Faithfulness: Our time-aware faithfulness metric examines whether details in the impact summary, i.e., the impact description elaborating on the dominant intent of an impact period, is grounded in the paper’s citation contexts (i.e., the LLM’s input context). We first split the impact summaries into impact descriptions representing different time periods (as defined in the output schema in Appendix C). Next, we instruct an LLM to verify each against citations from the corresponding period. Note that a summary impact description is verified against many citations at once and not in a simple pairwise entailment, since a single impact description can encompass information from multiple citations or discuss trends, which are only possible to verify by inspecting many sources (e.g., “the paper is frequently used for…”). The evaluator LLM is instructed to assess the impact description against the provided citations, determine its faithfulness, and justify the decision with a list of supporting citations. This procedure resembles the task of a human annotator evaluating the faithfulness of a machine-generated text Kim et al. (2024). (2) Coverage: Coverage is defined as the ratio of impact-revealing intents mentioned in the summary. For example, if a paper is cited for its method use, for inspiring new research, and for exposing limitations and refinement, the summary should reflect all these intents, including relevant details. Inspired by the evaluation rubrics for coverage proposed in Asai et al. (2024a), which defines it as the topics (themes) mentioned (i.e., the diversity of impact-revealing intents in our case), we develop a two-step evaluation. In the first step, we cluster the list of fine-grained intents under similar topics, e.g., “method used in the legal domain”. These labels might have slightly different granularity, which depends on how much details the members (i.e., intents) of that cluster offer. Next, the evaluator LLM takes both the impact summary and the list of cluster titles that this summary should ideally cover, and returns a list of topics that were actually covered. Finally, we divide the number of intents mentioned in the summary by the total number of impact-revealing intent clusters. Prompts for faithfulness and coverage are in Appendix D. (3) Citation Year Compliance: We implement this metric as a script that flags citations falling outside the target impact period.
Informativeness. We also assess the informativeness of impact summaries, as trustworthy ones may still be unhelpful if they lack meaningful insights. Grounded in our definitions (Section 2.1), the informativeness of an impact summary depends on its ability to clearly describe the paper’s direct influence on other papers, track the types of influence (intents) over time, and provide details inferred from the citation contexts. We define the following metrics: (1) Insightfulness measures how well the impact summary articulates the paper’s direct influence on other works. (2) Trend Awareness evaluates the extent to which the impact summary identifies and distinguishes between different periods of the paper’s influence over time. (3) Specificity assesses whether the impact summary includes concrete examples, such as techniques, influenced by the paper. To systematically evaluate these criteria, we leverage G-Eval Liu et al. (2023a), a framework that uses LLMs with chain-of-thought (CoT) reasoning to assess the impact summaries. By integrating structured evaluation steps, G-Eval enables LLMs to reason step-by-step, improving the reliability and depth of assessments. We transform our criteria into prompts in Appendix D.
To construct a new dataset for classifying citation contexts as “impact-revealing” or not, we start by building on the PST-Bench dataset Zhang et al. (2024a), which focuses on positive influence through “inspiration” and “motivation”. We select 1k impact-revealing citations from PST-Bench. Next, we augment this by first manually crafting textual patterns for both confirmation and correction intents, and use GPT-4 to generate variations. After crawling 200k random citation contexts (from the Semantic Scholar Academic Graph (S2AG) [^4] https://www.semanticscholar.org/product/api) and checking if they match any of the textual patterns, we randomly sample 1k instances of impact-revealing contexts (a total of 2k impact-revealing with the original selection from PST-Bench, covering both confirmatory and correct citations). Additionally, we sample 2k non-impact-revealing citations (“other”) directly from the PST-Bench dataset (references that were not annotated as influential). We also ensure none of these match any of our impact-revealing citation patterns. This results in a balanced 4k citation context dataset, which we release alongside this work. To validate the quality of our automatically collected impact-revealing examples, we manually reviewed a random sample of 100 instances and found that 90% were correctly labeled. The remaining 10% consisted of edge cases where surface patterns, such as the phrase “motivated by” (e.g., Voters are motivated by partisan social identities… (Greene 2004)), appeared to indicate impact but did not actually reflect the citation’s true intent. Details on the construction and textual patterns are in Appendix F.1.
3 Experiments
3.1 In-context-learning for fine-grained intent generation
Zero-shot vs. ICL. To evaluate LLMs’ ability to generate fine-grained intents and classify them as “impact-revealing” or “non-impact-revealing”, we manually select 10 citation contexts (Table 1 and Appendix E.1). Our selection criteria was based on diversity, a set that contains impact-revealing (praise, critique) and non-impact-revealing (“other”) intents. The “other” class includes incidental mentions or contexts lacking sufficient information to understand the reason behind it. We use our prompt (Appendix B) with and without these examples to generate and classify intents, using GPT4o-mini, of 200 randomly sampled citation contexts. Two annotators (a postdoc and a PhD student) assess: (i) whether the classification as impact-revealing is correct, (ii) whether the generated intent accurately reflects the citation reason, and (iii) whether the intent is concise, as LLM results can be highly verbose. Results, including per-field metrics and qualitative examples, are in Appendix E.2. We find that including examples significantly improves performance across all metrics, with gains of 29%, 55%, and 39% in precision, recall, and F1, respectively. Inter-annotator agreement shows a substantial Cohen’s kappa score of 0.68 on impact-revealing intent classification.
Effect of different number of shots (K). We find that performance improves significantly as the number of shots increases, plateauing around K = 50 with strong metrics (recall: 0.94, F1: 0.92); see Appendix F for details.
How does citation intent vary across fields? Table 2 presents an analysis of citation intents across 70k citation contexts from psychology, medicine, and computer science papers, with further breakdowns by recency and citation counts. Overall, psychology citations tend to lean towards impact-revealing, while computer science skews more toward non-impact-revealing (“other”) citations—except in more recent papers—and medicine shows a more balanced distribution. We observe that citations in psychology papers often exhibit a stronger subjective tone, especially in correction citations, such as: “some of those assumptions have been controversial” and “researchers disagree about whether the kinds of behaviors measured by particular implicit tests should be considered indicators of attitudes or something else”. In contrast, computer science shows a notable shift in recent years toward more impact-revealing citations, likely driven by the novelty and immediate influence of AI research before transitioning into the legacy phase, where citations become more background-oriented. This legacy effect is particularly evident in the highly cited subsets.
Cited Papers Citation Intents (%)
PS MD CS
All 6535 5347 3466
Recent 6535 5743 5842
Older 6535 5248 3367
Highly cited 6634 3664 3862
Less cited 6436 5644 3268
@@ 3.1 In-context-learning for fine-grained intent generation
Intent classifier Confirmatory Correction
Structural Scaffolds 0.83 0.50
Meaningful Citations 0.41 0.65
Multi-cite 0.81 0.58
Ours 0.88 0.98
Table 4: F1 scores for impact-revealing citations, broken down by confirmatory and correction categories.
@@ 3.2 Can existing intent classifiers detect impact-revealing citations?
Comparison with existing methods. We compare our classifier against the following popular intent classification methods: (i) Meaningful Citations Valenzuela et al. (2015): a supervised machine learning method to classify citation intents into meaningful or non-meaningful, by leveraging context and citation metadata; (ii) Structural Scaffolds Cohan et al. (2019): a multi-task model which enhances the main task by integrating auxiliary tasks; (iii) Multi-cite Lauscher et al. (2022): a multi-label classification method that predicts multiple intents using a fine-tuned SciBERT Beltagy et al. (2019). We map their intent classes to either “impact-revealing” or “other”. More details are in Appendix F.2. For our method, we use the ICL prompt, with K=50 and LLM=GPT4o-mini. For every test instance, we run our prompt 3 times (shuffling the order of the shots), and take a majority vote for the classification. Note that the results of the 3 runs have a full agreement of 72%. By full agreement, we mean all three runs predicted the same class, indicating stability to in-context example order.
Results. Numerical results are reported in Table 3. Our method demonstrates the best performance, with the most significant improvement in recall, outperforming the second-best method by 19 percentage points. This is particularly crucial for our goal in generating impact summaries from impact-revealing citations because the highest recall in this context means that our method ensures that fewer meaningful citations are missed. The qualitative results are shown in the appendix F.2. For a deeper analysis on the performance of these classifiers on confirmatory vs. correction citations, we report the F1 scores on these splits in Table 4. The table shows that our strong performance is largely driven by the intent classifier’s effectiveness at identifying impact-revealing citations with correction-related intents (e.g., method refinement or highlighting research gaps). To ensure a fairer comparison with prior methods, we restrict the results in Table 4 to citations from computer science papers, aligning with the domain used to train those models.